(p^4)^5/p^7 I think I got this one. Is it p^2 ?
Answer by checkley71(8403)

(p^4)^5/p^7=p^20/p^7=p^(20-7)=p^13
Raising a power to a power is a multiplication process.
Multiplying a power by a power is an addition process.